### v1

# Venkata Ramana Makkapati Honda Aircraft Company

Abstract
This paper introduces a novel reachability problem for the scenario where one agent follows another agent using the constant bearing pursuit strategy, and analyzes the geometry of the reachable set of the follower. Key theoretical results are derived, providing bounds for the associated dependent reachable set. Simulation results are presented to empirically establish the shape of the dependent reachable set. In the process, an original optimization problem for the constant bearing strategy is formulated and analyzed.
1 Introduction
Reachability theory is widely applied to address collision avoidance and safety in the planning and control of multi-agent systems [althoff2014online, malone2017hybrid, wang2020infusing]. The reachability analysis involves identifying the set that contains all the states an agent can reach given the initial conditions and the allowable control inputs, as a function of time. In instances where one agent follows another agent using a feedback strategy, our problem of interest is to identify the follower’s reachable set, which is referred to as the dependent reachable set (DRS) in this paper. The idea of a DRS was introduced by Comandur et al. to analyze deception relevance in games with asymmetric information [comandur2024desensitization]. The problem becomes relevant when the leader (the agent being followed) is a faulty robot [park2017fault], or when the leader itself is the opponent in strategic engagements [ho2022game].
The reachability problem, a dual of the controllability problem, was extensively studied for linear systems [gayek1991survey]. Computing reachable sets for nonlinear continuous-time systems is more complex, as an analytical expression for the set is not always guaranteed [althoff2021set]. Consequently, several approaches including those based on optimal control theory [mitchell2005time, chen2018hamilton], barrier functions [kong2018reachable], and trajectory sensitivity analysis [ramdani2008reachability, chuchu2016locally], were developed to estimate reachable sets. More recently, there has been an increased focus on reachability analysis for multi-agent systems [thapliyal2024algorithm, wang2021formal], and neural network controllers [gates2023scalable, everett2021reach].
The constant bearing pursuit strategy is also known as “parallel navigation” in the robotics literature [strydom2015biologically], and is popular among missile guidance problems [Shneydor], and differential games [isaacs1999differential]. The guidance strategy was applied to mobile robots to reach a moving target [belkhouche2007parallel, rafie2010time].
This paper characterizes DRS for an agent that employs the constant bearing strategy to pursue another agent. To the best of authors’ knowledge, the concept of DRS has not been studied, and is new to the literature. The contributions of this paper are as follows: a) the concept of DRS is introduced for multi-agent systems; b) the geometry of DRS is characterized for an agent that employs the constant bearing pursuit strategy; c) a novel optimization problem for scenarios involving the constant bearing strategy is formulated and discussed.
The paper is organized as follows. Section 2 presents the problem formulation and the underlying assumptions. Section 3 discusses the theory of reachable sets. Section 4 provides key results, both theoretical and simulation, that describe the geometry of DRS for the constant bearing strategy. Section 5 studies the optimization problem to find the closest and farthest points from the leader that the agent can reach while employing the constant bearing strategy. Section 6 concludes the paper and provides some directions for future work.
2 Problem Statement
Consider a scenario involving two agents (denoted using the subscripts $I$ and $D$) moving in the Cartesian plane at a constant speed by controlling their respective heading angles. The agents’ dynamics can be expressed as
$\dot{\mathbf{x}}_{i}(t)$ $=\mathbf{u}_{i}(t),\quad\mathbf{x}_{i}(0)=\mathbf{x}_{i}^{0},\quad\|\mathbf{u}_{i}(t)\|_{2}=v_{i},$ (1)
[…]
Without loss of generality, we consider $\mathbf{x}^{0}_{D}$ to be the origin, and $\mathbf{x}^{0}_{I}=[a,0]^{\top}$. Consequently, the initial line of sight (LOS) is parallel to the horizontal axis. Since the dependent agent follows the constant bearing pursuit strategy, it chooses a heading angle such that the LOS does not rotate, and the relative distance from the independent agent is reduced at a given time instant [makkapati2018optimal]. Therefore,
$\psi_{D}(t)=\sin^{-1}\left(\dfrac{v_{I}}{v_{D}}\sin\psi_{I}(t)\right).$ (2)
In the remainder of this paper, we consider only the case where $\mathbf{u}_{D}(t)=[v_{D}\cos\psi_{D}(t),~v_{D}\sin\psi_{D}(t)]^{\top}$ for $\psi_{D}(t)$ given in (2)
Assumption 2.1.
The speed of the dependent agent is greater than that of the independent agent, i.e. $v_{D}>v_{I}$.
Note that $v_{D}\geq v_{I}$ is a necessary and sufficient condition for $\psi_{D}(t)$ to exist, given any $\psi_{I}(t)\in(-\pi,\pi]$, as can be inferred from (2). Furthermore, given $v_{D}\geq v_{I}$ and $\psi_{I}(t)\in(-\pi,\pi]$, $\psi_{D}(t)\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$. In this paper, $v_{D}>v_{I}$ is considered for presenting theoretical and empirical results. The analysis corresponding to the limiting case of $v_{D}=v_{I}$ is presented in Section 4.1.
Given the dynamics of the independent agent per (1), its trajectory $\mathbf{x}_{I}(t),~t\geq 0$ is a function of the control function $\mathbf{u}_{I}\in\mathcal{U}_{I}$. For every trajectory of the independent agent, there is a corresponding trajectory of the dependent agent that is obtained using the constant bearing feedback strategy. Now, the problem statement examined in this paper is presented below.
Problem 2.2.
For the set of all feasible trajectories of the independent agent, determine the corresponding set, $\mathcal{D}(t)$, of positions that will be reached by the dependent agent.
In other words, given that the dependent agent is committed to a constant bearing pursuit strategy, what is the set of points that the independent agent can force the dependent agent to reach? When $v_{D}>v_{I}$, every possible trajectory of the independent agent will result in a capture in finite time by the dependent agent that follows the constant bearing pursuit strategy. Here, capture refers to the event $\|\mathbf{x}_{D}(t)-\mathbf{x}_{I}(t)\|_{2}\approx 0$. In this regard, the paper considers that the set $\mathcal{D}(t)$ comprises only those points that correspond to the active pursuit trajectories at a given time. By active pursuit trajectories, we mean those trajectories of the dependent agent that are yet to result in capture for the corresponding trajectory of the independent agent. The maximum capture time $t_{c}=a/(v_{D}-v_{I})$ corresponds to an instance when the independent agent chooses the control input $\mathbf{u}_{I}(t)=[v_{I},0]^{\top}$ for all time until capture. In the pursuit-evasion literature, this chosen control input of the independent agent corresponds to what is known as pure evasion, which guarantees the maximum capture time. Note that as $v_{I}\rightarrow v_{D}$, $t_{c}\rightarrow\infty$.
In this paper, a point from the 2D plane is either declared as a column vector of the form $\mathbf{x}=[x,y]^{\top}$, or denoted in the Cartesian form $X\left(x,y\right)$ as an ordered pair. The following section presents important results for the reachable set of the independent agent, which will be used to characterize the set $\mathcal{D}(t)$ for the dependent agent.
3 Reachable Sets
A standard definition for the independent agent’s reachable set at time $t\geq 0$ is given below.
Definition 3.1.
The independent agent’s reachable set, $\mathcal{R}_{I}(\mathbf{x}_{I}^{0},t)$, at time $t\geq 0$ with the initial state at $\mathbf{x}_{I}^{0}$ is the set of all points that the agent can reach at time $t$:
[…]
A closed-form expression for the reachable set of the independent agent, per the dynamics in (1), can be obtained using the following lemma. Hereafter, the dependency of the reachable set on the initial condition, which is apparent, is dropped for brevity, and the reachable set at time $t$ is denoted as $\mathcal{R}_{I}(t)$.
Lemma 3.2.
The reachable set of an agent in a one-dimensional environment with the dynamics
$\dot{x}(t)=w(t),\quad x(0)=0,$ (4)
at time $t=T\in[0,\infty)$ is [T,T][-T,T], where $x(t)\in\mathbb{R}$, the control function $w\in\mathcal{W}$, and $\mathcal{W}$ is the set of all piecewise continuous functions in time $t\in[0,\infty)$ with range $\{-1,1\}$.
Proof.
Consider the single-switching control function
$w_{s}(t)=\begin{cases}1,\quad\text{if }0\leq t<t_{s},\\ -1,\quad\text{if }t_{s}\leq t\leq T,\end{cases}$ (5)
where $t_{s}\in(0,T)$ is the switching time. Any point $x_{s}\in(-T,T)$ can be reached with the control function in (5) using switching time
$t_{s}$ $=\dfrac{x_{s}+T}{2}.$ (6)
It can be observed that the switching times vary between 0 and $T$ to span all the points in the set (T,T)(-T,T). Since the agent moves at a unit speed, the maximum distance it can reach from its initial point at time $T$ is $T$. Therefore, in the one-dimensional case, the reachable set at time $T$ is [T,T][-T,T]. ∎
Lemma 3.3.
The reachable set of the independent agent, per the dynamics in (1), at time $T\in[0,\infty)$ is the circle with its center at $\mathbf{x}_{I}^{0}$ and radius $v_{I}T$.
Proof.
Without loss of generality, let $\mathbf{x}_{I}^{0}$ be the origin. Using Lemma 3.2, all points on the line segment joining the origin and the point $v_{I}T\cos\theta,v_{I}T\sin\theta]^{\top}$ can be reached with the control function of the form
$\mathbf{u}_{1s}(t)=\begin{cases}[v_{I}\cos\theta,v_{I}\sin\theta]^{\top},\quad&\text{if }0\leq t<t_{s},\\ [-v_{I}\cos\theta,-v_{I}\sin\theta]^{\top},\quad&\text{if }t_{s}\leq t\leq T,\end{cases},$ (7)
using an appropriate switching time, for all $\theta\in(-\pi/2,\pi/2]$. The points on the boundary of the reachable set, which are of the form $v_{I}T\cos\theta,v_{I}T\sin\theta]^{\top}$, are reached by choosing the control $\mathbf{u}_{I}(t)=[v_{I}\cos\theta,v_{I}\sin\theta]^{\top}$, $t\in[0,T]$. Therefore, in the two-dimensional case
[…]
4 Dependent Reachable Sets
The solution to Problem 2.2 involves identifying a specialized reachable set of the dependent agent, termed the dependent reachable set (DRS). DRS is defined for the agent that determines its control input using a feedback control function of the form $\mathbf{u}_{D}(t)=f(\mathbf{x}_{D}(t),\mathbf{x}_{I}(t),\mathbf{u}_{I}(t))$, which is a function of the instantaneous state and control input of the independent agent. A formal definition for the DRS is presented below.
Definition 4.1.
The dependent reachable set of the dependent agent that employs the feedback strategy $f$, at time $t\geq 0$ is defined as
[…]
$\mathbf{x}=\mathbf{x}_{D}^{0}+\int_{0}^{t}f(\hat{\mathbf{x}}_{D}(\tau),\hat{\mathbf{x}}_{I}(\tau),\hat{\mathbf{u}}_{I}(\tau))\mathrm{d}\tau\Bigg\},$ (9)
where $\hat{\mathbf{x}}_{D}(\tau)=\mathbf{x}_{D}^{0}+\int_{0}^{\tau}f(\hat{\mathbf{x}}_{D}(\eta),\hat{\mathbf{x}}_{I}(\eta),\hat{\mathbf{u}}_{I}(\eta))\mathrm{d}\eta$ and $\hat{\mathbf{x}}_{I}(\tau)=\mathbf{x}^{0}_{I}+\int_{0}^{\tau}\hat{\mathbf{u}}_{I}(\eta)\mathrm{d}\eta$.
In this paper, we analyze dependent reachable sets in the case where the dependent agent follows the constant bearing strategy. We denote the DRS by $\mathcal{D}(t)$, dropping the dependencies on the initial conditions $\mathbf{x}_{D}^{0}$, $\mathbf{x}_{I}^{0}$ for the sake of notation brevity.
For the problem described in Section 2, the reachable set of the independent agent $\mathcal{R}_{I}(t)$ is a circular region with its center at $\left(a,0\right)$ and radius $v_{I}t$ (see Lemma 3.3). Similarly, the reachable set of the dependent agent $\mathcal{R}_{D}(t)$ is also a circle with its center at the origin and radius $v_{D}t$:
[…]
The boundary of the dependent agent’s reachable set is denoted by $\partial\mathcal{R}_{D}(t)$. Note that
$\mathcal{D}(t)\subseteq\mathcal{R}_{D}(t).$ (11)
Figure 1: Schematics of the two distinct scenarios that are considered to characterize the dependent reachable set for the constant bearing strategy
4.1 Theoretical Results
This subsection presents the characterization of the DRS for the constant bearing strategy at time $t$, including the points $P_{1}\left(t\sqrt{v_{D}^{2}-v_{I}^{2}},v_{I}t\right)$ and $P_{2}\left(t\sqrt{v_{D}^{2}-v_{I}^{2}},-v_{I}t\right)$ lying on $\partial\mathcal{R}_{D}(t)$, which are obtained as a result of the constant bearing strategy attributes.
As shown in Fig. 1, two different scenarios are considered to develop proofs that provide the shape of the DRS at a given instant in time $t\in[0,t_{c}]$. The vertical diameter of $\mathcal{R}_{I}(t)$ (depicted using a dashed line in Fig. 1) is the line joining $(a,v_{I}t)$ and $(a,-v_{I}t)$, which is perpendicular to the initial LOS (the horizontal axis). The first scenario in Fig. 1 depicts the case where the vertical diameter of $\mathcal{R}_{I}(t)$ does not completely lie in the interior of the circle $\mathcal{R}_{D}(t)$. The first scenario could correspond to the circle $\mathcal{R}_{D}(t)$ either intersecting or not intersecting the vertical diameter of $\mathcal{R}_{I}(t)$. The earliest time instant when $\partial\mathcal{R}_{D}(t)$ first intersects the vertical diameter is at $t_{1}=a/v_{D}$. Note that at time $t_{1}$, the initial position of the independent agent is part of the boundary of the dependent agent’s reachable set. At the same time, it can be deduced that the latest time instant when $\partial\mathcal{R}_{D}(t)$ intersects the vertical diameter is at $t_{2}=a/\sqrt{v_{D}^{2}-v_{I}^{2}}$. In summary, the first scenario in Fig. 1 corresponds to instances occurring in the time interval $0\leq t<t_{2}$.
Consequently, the second scenario in Fig. 1 depicts the case where $\partial\mathcal{R}_{D}(t)$ fully encompasses the vertical diameter of $\mathcal{R}_{I}(t)$. The instances corresponding to the second scenario occur in the time interval $t_{2}<t\leq t_{c}$. The following lemmas help characterize the DRS for the constant bearing strategy.
Figure 2: Characterization of the DRS for $0\leq t\leq t_{2}$
Lemma 4.2.
[…]
Proof.
For the constant bearing strategy, the LOS, which is initially aligned along the horizontal axis, does not rotate until capture. As a result, $y_{D}(t)=y_{I}(t)$ for all time $t\in[0,t_{c}]$, and the vertical coordinate of the dependent agent is bounded by $\pm v_{I}t$. Therefore,
[…]
Lemma 4.3.
[…]
Proof.
Given $\psi_{D}(t)$ is the heading angle of the dependent agent at time $t$, per (2),
$\underset{\psi_{I}(t)\in(-\pi,\pi]}{\arg\min}~\cos\psi_{D}(t)=\pm\dfrac{\pi}{2}.$ (13)
Consequently, at time $t$, the minimum horizontal speed of the dependent agent is
$\underset{\psi_{I}(t)\in(-\pi,\pi]}{\min}~v_{D}\cos\psi_{D}(t)=\sqrt{v_{D}^{2}-v_{I}^{2}}.$ (14)
As a result, the minimum horizontal distance that the dependent agent can travel within time $t$ is $t\sqrt{v_{D}^{2}-v_{I}^{2}}$. Note that the dependent agent always has a non-negative horizontal speed since $\psi_{D}(t)\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$, for $t\in[0,t_{c}]$ and $v_{D}\geq v_{I}$. Therefore,
[…]
From Lemmas 4.2 and 4.3, it can be deduced that $\mathcal{D}(t)$ will lie in the intersection of the regions given in (10), (12), and (15). The boundaries of these three regions intersect at the points $P_{1}$ and $P_{2}$. The intersection region for the first scenario ($0\leq t\leq t_{2}$) is shown in Fig. 2. Note that at time $t_{2}$, the line joining the points $P_{1}$ and $P_{2}$ (which lie on $\partial\mathcal{R}_{D}(t)$) coincides with the vertical diameter of $\mathcal{R}_{I}(t)$. The following theorem establishes that the DRS for the first scenario is the intersection region itself.
Theorem 4.4.
For $0\leq t\leq t_{2}$,
[…]
Proof.
Expressions in (10) and (15) establish that the sets $\|\mathbf{x}\|_{2}\leq v_{D}t$ and $x\geq t\sqrt{v_{D}^{2}-v_{I}^{2}}$ bound the DRS for all time $t\in[0,t_{c}]$. The proof involves showing that for every point $\mathbf{x}$ from the intersection of the aforementioned sets, there exists a control function $\mathbf{u}_{D}(.)\in\mathcal{U}_{D}$ which drives the dependent agent following the constant bearing strategy to that $\mathbf{x}$. In this regard, for time $t\in(0,t_{2}]$ and $\theta\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$, consider the switching control function of the type
$\mathbf{u}_{2s}(\tau)=\begin{cases}[v_{I}\cos\theta,v_{I}\sin\theta]^{\top},\quad&\text{if }0\leq\tau<t_{s},\\ \dfrac{\pi}{2},\quad&\text{if }t_{s}\leq\tau<\dfrac{t+t_{s}}{2},\\ -\dfrac{\pi}{2},\quad&\text{if }\dfrac{t+t_{s}}{2}\leq\tau\leq t,\end{cases},$ (17)
for the independent agent. Figure 3 indicates the resulting path (lines with arrows depicting the direction of motion) of the independent agent for the control function in (17). The corresponding horizontal coordinate of the dependent agent can be obtained as
$t_{s}\sqrt{v_{D}^{2}-v_{I}^{2}\sin^{2}\theta}+(t-t_{s})\sqrt{v_{D}^{2}-v_{I}^{2}}.$ (18)
The corresponding vertical coordinate for both agents can be obtained as $t_{s}v_{I}\sin\theta$.
For a constant vertical coordinate $y=t_{s}v_{I}\sin\theta$, $t_{s}\in[y/v_{I},t]$, the dependent agent’s horizontal coordinate in (18) can be rewritten as
$\sqrt{t^{2}_{s}v_{D}^{2}-y^{2}}+(t-t_{s})\sqrt{v_{D}^{2}-v_{I}^{2}}.$ (19)
From (19), it can be observed that all the points from the intersection region in (16) for a given $y\in[-v_{I}t,v_{I}t]$ are spanned by $t_{s}\in[y/v_{I},t]$, and correspond to the points reached by the independent agent using the control function in (17) for $\pi/2\leq\theta\leq\theta_{l}$ (when $y\geq 0$) or $\theta_{l}\leq\theta\leq-\pi/2$ (when $y<0$). Note that $\theta_{l}=\tan^{-1}(y/v_{I}t)$
Figure 3: Independent agent’s path for the control function in (17), and the resulting points of both agents for a given vertical coordinate $y$.
Next, we show that all the resulting trajectories of the dependent agent corresponding to the control function in (17), and $\theta\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$ will not result in capture for $0\leq t<t_{2}$. For a capture to occur at time $t$ with the independent agent’s control function in (17), the horizontal coordinates of both agents should be the same. Consequently,
$t_{s}\sqrt{v_{D}^{2}-v_{I}^{2}\sin^{2}\theta}+(t-t_{s})\sqrt{v_{D}^{2}-v_{I}^{2}}=a+v_{I}t_{s}\cos\theta,$
$\implies t=\dfrac{a}{\sqrt{v_{D}^{2}-v_{I}^{2}}}+\dfrac{t_{s}}{\sqrt{v_{D}^{2}-v_{I}^{2}}}\Bigg[v_{I}\cos\theta$
$~~~~~~~~~~~~~~+\sqrt{v_{D}^{2}-v_{I}^{2}}-\sqrt{v_{D}^{2}-v_{I}^{2}\sin^{2}\theta}\Bigg]$ (20)
Since $t_{s}\geq 0$, $v_{D}>v_{I}$ and $\theta\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$, (20) yields the capture time $t\geq a/(\sqrt{v_{D}^{2}-v_{I}^{2}})=t_{2}$. ∎
As discussed in [makkapati2019optimal], the constant bearing strategy and the Apollonius circle have close connections. The Apollonius circle is defined for a pursuer-evader pair as the set of capture points when the pursuer follows the constant bearing strategy, and the evader chooses a constant heading angle. The geometric properties of the Apollonius circle were discussed in [ramana2017pursuit]. The center and radius of the Apolloinus circle can be obtained as $\left(av_{D}^{2}/(v^{2}_{D}-v_{I}^{2}),0\right)$ and $av_{D}v_{I}/v_{D}^{2}-v_{I}^{2}$, respectively.
It has been proven that the line segment joining the two tangent points $A_{1}$ and $A_{2}$, in Fig. 2, passes through the initial position of the independent agent [ramana2017pursuit]. Furthermore, the line joining $A_{1}$ and $A_{2}$ subtends an angle of $2\sin^{-1}(v_{I}/v_{D})$ at the initial position of the dependent agent, and the line is perpendicular to the initial LOS. At the same time, it is also evident that the line joining the points $P_{1}$ and $P_{2}$ subtends an angle of $2\sin^{-1}(v_{I}/v_{D})$ at the initial position of the dependent agent, and the line is perpendicular to the LOS. Therefore, the points $A_{1}$, $P_{1}$, and the initial position of the dependent agent are collinear. Similarly, the points $A_{2}$, $P_{2}$, and the initial position of the dependent agent are collinear. It can be noted that at time $t_{2}$, the initial position of the independent agent $x^{0}_{I}$ lies on the line segment $\overline{P_{1}P_{2}}$. Consequently, $\overline{P_{1}P_{2}}$ coincides with line segment $\overline{A_{1}A_{2}}$ at time $t_{2}$.
Corollary 4.5.
[…]
Proof.
The proof directly follows from Lemmas 4.2, 4.3. ∎
The above corollary provides a bound for the DRS in the second scenario ($t_{2}<t\leq t_{c}$). The bound is characterized by the points $P_{1}$ and $P_{2}$, and is depicted in Fig. 4 using a bold line. It is important to note that the above corollary, while providing a tight bound for the DRS, does not define the boundary of the DRS. The mathematical proof to establish the boundary of the DRS for $t_{2}<t\leq t_{c}$ is elusive. While a formal mathematical proof remains an open challenge, we present compelling evidence using simulation results (in the following subsection) that supports our hypothesis below.
Hypothesis 4.6.
For $t_{2}<t\leq t_{c}$,
[…]
$x\geq\dfrac{a^{2}+v_{D}^{2}t^{2}-v_{I}^{2}t^{2}}{2a}\Bigg\}.$ (21)
The vertical line $x=(a^{2}+v_{D}^{2}t^{2}-v_{I}^{2}t^{2})/2a$ contains the points $Q_{1}$ and $Q_{2}$, which are the intersection points for the boundaries of the reachable sets of both agents ($\partial\mathcal{R}_{I}$ and $\partial\mathcal{R}_{D}$), as shown in Fig. 4. In the following subsection, it is empirically shown that the DRS for time $t_{2}<t\leq t_{c}$ is the shaded region in Fig. 4, which is characterized by the minor segment $Q_{1}P_{x}Q_{2}$.
Figure 4: Characterization of the DRS for $t_{2}<t\leq t_{c}$
The Limiting Case of $v_{D}=v_{I}$
In the case of $v_{D}=v_{I}$, at time $t$, the points $P_{1}$ and $P_{2}$ are obtained as $(0,v_{I}t)$ and $(0,-v_{I}t)$, respectively. Therefore, the points $P_{1}$, $P_{2}$, and the initial point of the dependent agent are collinear. The line joining the three points forms the vertical diameter of the circle $\partial\mathcal{R}_{D}(t)$ (boundary of the dependent agent’s reachable set), which can be visualized in Fig. 5
Note that as $v_{I}\rightarrow v_{D}$, $t_{2}\rightarrow\infty$. Consequently, only the first scenario discussed in Section 4.1 is relevant for the limiting case, and the second scenario is irrelevant. Furthermore, it can be observed that the conclusion drawn in the proof of Theorem 4.4 is valid for the case of $v_{D}=v_{I}$. Consequently, the DRS is given by
[…]
Figure 5: DRS for the limiting case $v_{D}=v_{I}$
(a) $t=0.6\leq t_{1}$
(b) $t=1.0=t_{1}$
(c) $t=1.2\approx t_{2}$
(d) $t_{2}<t=1.4<t_{c}$
(e) $t_{2}<t=1.8<t_{c}$
(f) $t=2.0=t_{c}$
Figure 6: Point cloud-based simulation results depicting the evolution of DRS for $0<t\leq t_{c}$
4.2 Simulation Results
This subsection presents simulation results based on discrete-time point-cloud propagation to observe the evolution of the dependent reachable sets for the time $0\leq t\leq t_{c}$. The initial positions of both agents are $x_{D}^{0}(0,0)$ and $x_{I}^{0}(1,0)$ and their speeds are set to $v_{D}=1$ and $v_{I}=0.5$ respectively. Simulations are carried out with the time-step $\Delta t=0.2$ until all the independent agents are captured. To generate a point cloud, we consider that at a given time instant, an independent agent chooses eighteen equally-spaced heading angles from $(-\pi,\pi]$. Consequently, at the $n^{th}$ time step, $18^{n}$ points are generated for the independent agent. At any time instant, every independent agent point has an associated dependent agent point obtained via the constant bearing strategy. For the simulation parameters mentioned above, $t_{1}=1.0$, $t_{2}\approx 1.155$ and $t_{c}=2.0s$.
Figure 6 presents the results obtained from the point cloud-based simulations. The blue color corresponds to the dependent agent, and the red color corresponds to the independent agent. The dashed circles indicate the boundaries of the agents’ reachable sets. The green line indicates the boundary of the DRS. The inverted triangles represent the initial positions of the two agents. The points in the red point cloud represent uncaptured independent agents, and the points in the blue point cloud represent the active dependent agents that are yet to capture their corresponding independent agents (refer to Section 2). As the simulation progresses, when a dependent agent captures an independent agent, that particular pair is eliminated from their respective point clouds. Figures 6(a) and 6(b) represent two instances of the scenario $0<t\leq t_{1}$, and Fig. 6(c) represents the case where $t\approx t_{2}$. Figures 6(d)-6(f) depict three instances of $t_{2}<t\leq t_{c}$.
[…]
In Figs. 6(d) and 6(e), first, it can be observed that the minor segment $P_{1}P_{x}P_{2}$ contains points that are not part of the active region of the independent agent’s reachable set. Note that in the case of constant bearing, the relative distance between the agents reduces along the LOS. Therefore, the points in the minor segment $P_{1}P_{x}P_{2}$ that are not part of the independent agent’s reachable set cannot be part of the DRS. Furthermore, there are no blue points in the region between the line segments $\overline{P_{1}P_{2}}$ and $\overline{Q_{1}Q_{2}}$. It can be observed that the chord $\overline{Q_{1}Q_{2}}$ is part of the boundary of the blue point cloud when $t_{2}<t\leq t_{c}$, and the minor segment $Q_{1}P_{x}Q_{2}$ contains all the blue points when $t_{2}<t\leq t_{c}$. A more enhanced image of the interaction between the reachable sets is shown in the upper left corner of each figure. Simulations have been conducted for different parameters ($v_{D}$, $v_{I}$, etc.), which have not been included for brevity, and the results provided the same observations. Thus, it could be empirically stated that the DRS, initially represented by the segment $P_{1}P_{x}P_{2}$, reaches a maximum area at $t=t_{2}$ and subsequently shrinks to the segment $Q_{1}P_{x}Q_{2}$, which is contained within the segment $P_{1}P_{x}P_{2}$. Figure 6(f) ($t=t_{c}$) presents the case where the dependent agent’s reachable set fully encompasses the independent agent’s reachable set, and all the points from both red and blue point clouds ended in capture.
5 The Optimization Problem
This section formulates an optimization problem and provides empirical results pertaining to the constant bearing strategy, and the corresponding DRS. An approach to identifying the DRS is to analytically define the set of all the points that the dependent agent can reach using the constant bearing strategy for a given point from the independent agent’s reachable set. In this regard, consider a point $\mathbf{x}_{I}(t)=\mathbf{x}=[x,y]^{\top}\in\mathcal{R}_{I}(t)$, which the independent agent reaches at time $0\leq t\leq t_{c}$. Note that if $\|\mathbf{x}\|_{2}<v_{I}t$, there are infinitely many trajectories that the independent agent can take to reach the point $\mathbf{x}$. The goal is to find all the points that the dependent agent, which follows the constant bearing strategy, can reach for the given $\mathbf{x}$.
As formulated in Section 2, since the dependent agent follows the constant bearing strategy and given the initial conditions of both agents (origin and $(a,0)$), the vertical coordinates of both agents are the same for all time $t\geq 0$. Therefore, the vertical coordinate of the dependent agent at time $t$ is the same as the independent agent’s vertical coordinate, which is $y$. Consequently, we intend to find the minimum and the maximum horizontal coordinates that the dependent agent can reach at the time instant $t$. Since $\dot{y}_{D}(\tau)=\dot{y}_{I}(\tau)$ in this scenario, $\dot{x}_{D}(\tau)=\sqrt{v_{D}^{2}-\dot{y}_{I}^{2}(\tau)}$. In this regard, consider the problem of obtaining the extrema for the functional
$\int_{0}^{t}\sqrt{v_{D}^{2}-\dot{y}_{I}^{2}(\tau)}~\mathrm{d}\tau,$ (23)
for the constraints
$\dot{x}_{I}^{2}(\tau)+\dot{y}_{I}^{2}(\tau)=v_{I}^{2},$ (24)
$\mathbf{x}_{I}(0)=[a,0]^{\top},\quad\mathbf{x}_{I}(t)=\mathbf{x}=[x,y]^{\top}.$ (25)
The constraint in (24) can be absorbed into the optimization by considering the Lagrangian
$L=\sqrt{v_{D}^{2}-\dot{y}_{I}^{2}(\tau)}+\lambda(\tau)\left[\dot{x}_{I}^{2}(\tau)+\dot{y}_{I}^{2}(\tau)-v_{I}^{2}\right],$ (26)
where $\lambda(\tau)$ is the auxiliary variable. The first-order necessary conditions for the extremum points can be obtained using the corresponding Euler-Lagrange equations given below.
$\dfrac{\partial L}{\partial x_{I}}-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(\dfrac{\partial L}{\partial\dot{x}_{I}}\right)=-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(2\dot{x}_{I}(\tau)\lambda(\tau)\right)=0,$ (27)
$\dfrac{\partial L}{\partial y_{I}}-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(\dfrac{\partial L}{\partial\dot{y}_{I}}\right)=-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\Bigg(-\dfrac{\dot{y}_{I}(\tau)}{\sqrt{v_{D}^{2}-\dot{y}_{I}^{2}(\tau)}}$
$~~~~~~~~~~~~~~~~~~~~~~~~~+2\lambda(\tau)\dot{y}_{I}(\tau)\Bigg)=0.$ (28)
Consequently,
$\dot{x}_{I}(\tau)\lambda(\tau)=c_{1},$ (29)
$\dot{y}_{I}(\tau)\left[1/(\sqrt{v_{D}^{2}-\dot{y}_{I}^{2}(\tau)})-2\lambda(\tau)\right]=c_{2},$ (30)
for $0\leq\tau\leq t$, where $c_{1}$ and $c_{2}$ are constants.
Equations (27) and (28) cannot be solved further to obtain analytical expressions of the extrema. Therefore, simulation-based empirical studies were conducted to understand the nature of the extrema in the case where the independent agent follows a single-switching control function to reach the point $\mathbf{x}$. Note that in such cases, the locus of points at which the independent agent switches between constant control inputs is the ellipse
$\mathcal{E}=\Big\{\mathbf{x}_{s}=[x_{s},y_{s}]^{\top}:\|\mathbf{x}_{s}-\mathbf{x}_{I}(0)\|+\|\mathbf{x}_{s}-\mathbf{x}\|=v_{I}t\Big\}.$ (31)
The following hypotheses are proposed on the basis of empirical evidence.
Hypothesis 5.1.
The extremum points for the optimization problem in (23)-(25) correspond to the independent agent’s trajectories with a single-switching control function such that $\mathrm{abs}(\dot{x}(\tau))$ and $\mathrm{abs}(\dot{y}(\tau))$ are constants for all time $0\leq\tau\leq t$.
Hypothesis 5.2.
Per Hypothesis 5.1, the points in the Cartesian plane where the independent agent switches the control input for the maxima of the optimization problem in (23)-(25) are given by $\underset{\mathbf{x}_{s}\in\mathcal{E}}{\max}~x_{s}$ and $\underset{\mathbf{x}_{s}\in\mathcal{E}}{\min}~x_{s}$. The points for the minima are given by $\underset{\mathbf{x}_{s}\in\mathcal{E}}{\max}~y_{s}$ and $\underset{\mathbf{x}_{s}\in\mathcal{E}}{\min}~y_{s}$.
The above hypotheses are demonstrated using simulation results. Figure 7 depicts four cases from the numerous simulations that were performed to analyze the extrema of the optimization problem in (23)-(25). The simulations were conducted with $\mathbf{x}_{I}(0)=[0,0]^{\top}$ (red square), $v_{D}=1$, $v_{I}=0.5$, $t=8$. The four points (blue, black, purple and green coloured inverted triangles) chosen within $\mathcal{R}_{I}(t)$ (interior of the red circle) represent a subset of the independent agent’s possible locations at time $t$. In each case, the corresponding coloured ellipse represents the locus of switch points $\mathcal{E}$, as given in (31), that the independent agent could execute with a single-switching control function to reach the selected point from $x_{I}(0)$. For each ellipse, the integral in (23) is evaluated for 360 equally spaced points (in terms of angular displacement) on the ellipse to identify the maximum and minimum points. On each ellipse in Fig. 7, the maxima points are indicated using filled circles, and the minima points are represented using empty circles. Hence, it can be observed that the maxima points always correspond to the points with the maximum and the minimum horizontal coordinates on the ellipse. Similarly, the minima points coincide with the points having the maximum and the minimum vertical coordinates on the ellipse. As $\mathbf{x}_{I}(t)$ varies between the origin and the boundary of $\mathcal{R}_{I}(t)$, the locus of the switch points changes from circular to elliptical with increasing eccentricity.
Figure 7: Simulation results that empirically substantiate Hypotheses 5.1 and 5.2.
6 Conclusion
This paper analyzes the “reachable set” of an agent (termed the dependent agent) in instances where it follows another agent (termed the independent agent) using the constant bearing feedback strategy. It is assumed that the speed of the dependent agent is greater than the speed of the independent agent. Theoretical results are presented that analytically characterize the dependent reachable set (DRS) for the instances where the dependent agent’s reachable set does not fully engulf the diameter of the independent agent’s reachable set. It is shown that the theoretical results can be extended to the limiting case where the speeds of both agents are equal. In instances where the dependent agent’s reachable set entirely engulfs either the diameter of the independent agent’s reachable set or the independent agent’s reachable set itself, a tight bound for the DRS is provided along with a hypothesis supported by simulation-based empirical evidence. The study of the DRS for the constant bearing strategy led us to examine a novel optimization problem, which resulted in more empirical evidence-based hypotheses related to the geometry of the constant bearing strategy and the ellipse.